# Dynamical Similarity in Multisymplectic Field Theory

### v3

@@ V Field Theory Classification
All theories we shall consider throughout will possess a metric; however, the precise status of this metric within the particular field theory of interest will affect the underlying multisymplectic bundle structure. In the simplest of cases, the metric is a fixed object, so that it neither evolves dynamically, nor may it assume different values. Whenever we consider theories defined on a flat static background, the metric is fixed, and will most often be the Minkowski metric $\eta_{\mu\nu}$.
Another possibility is that the metric enters the theory parametrically; in this case, it is not constrained to be fixed, and we may choose to assign different values to the metric, as the context befits. Whilst this kind of theory does allow the metric to be changed, the equations of motion are derived with the metric held fixed. For this reason, we must include the components of the metric as additional configuration space variables; however, since these are not variational in nature, the Lagrangian should not contain any metric velocity variables. Consequently, when treating theories of this type, we modify the configuration space $E$ and the corresponding jet bundle $J^{1}E$ in the following way
@@
in which $\textrm{Sym}_{2}^{d-1,1}(M)$ denotes the space of symmetric covariant tensors of rank two and Lorentzian signature (d1,1)(d-1,1) on $M$; sections of $\textrm{Sym}_{2}^{d-1,1}(M)$ then correspond to metrics on $M$. Note that we do not consider the jet bundle over the whole space $E\times_{M}\textrm{Sym}_{2}^{d-1,1}(M)$, but rather append the factor of $\textrm{Sym}_{2}^{d-1,1}(M)$ to the original jet bundle; this is precisely so that the components $g_{\mu\nu}$ of the metric are amongst the field variables, but the derivatives $g_{\mu\nu,\rho}$ are not.
The final possibility we shall consider is that of a variational metric; typically, theories which possess variational metrics are coupled to gravity, and are thus described by Lagrangians which do incorporate derivatives of the metric. As an example, we consider a principal $G$ bundle $\pi:P\rightarrow M$, with gauge group $G=U(1)$; denote by $p:C(P)\rightarrow M$ the bundle of connections of $P$, which we may identify with the quotient manifold $C(P)\cong J^{1}P/G$ [91]. Coupling this $U(1)$ gauge theory to gravity requires us to work with the following covariant configuration space
@@
upon which $(x^{\mu},A_{\mu};g_{\mu\nu})$ provide local coordinates. For concreteness, we have here restricted ourselves to a $U(1)$ gauge theory in $d=3+1$ spacetime dimensions. The gauge-invariant Lagrangian is now taken to be a function on
@@
Note that this space differs from that which we use when describing theories whose metrics are parametric; the use of the jet bundle $J^{1}(\textrm{Sym}_{2}^{3,1}(M))$ introduces velocity variables for the components $g_{\mu\nu}$, with the result that local coordinates on $\mathcal{P}$ are given by $(x^{\mu},A_{\mu},g_{\mu\nu},A_{\mu,\nu},g_{\mu\nu,\rho})$; when evaluated on the jet prolongation $j^{1}\phi:M\rightarrow\mathcal{P}$ of some section $\phi\in\Gamma(M,E)$, the velocity coordinates become the familiar derivatives $\partial_{\nu}A_{\mu}$ and $\partial_{\rho}g_{\mu\nu}$.
@@
As an illustration of the formalism we have developed thus far, let us consider the example of $N$ non-interacting real scalar fields $\phi^{a}$ (with $a=1,\,\cdots,N$) moving in some curved, $d$-dimensional spacetime $M$, equipped with a Lorentzian metric $g_{\mu\nu}$ of signature [^8] Our use of the mostly-minus convention is primarily so that we may retain a positive kinetic term; when studying General Relativity, we shall revert to the more customary (,+,+,)(-$\cdots$) signature. $(+,-,-,\cdots)$. Whilst we allow for a background different from Minkowski spacetime, we shall impose that the metric be non-variational, and thus parametric in nature. In accordance with the discussion of the previous section, this requires us to make the replacements
@@
where $E$ is the trivial bundle $E=M\times\mathbb{R}^{N}\rightarrow M$, which would be an appropriate covariant configuration space for a flat background. Consequently, local coordinates on $\mathcal{C}$ are $(x^{\mu},\phi^{a},\phi^{a}_{\mu};g_{\mu\nu})$, where we have used a semicolon to separate the variational and parametric degrees of freedom. The Lagrangian function $L:\mathcal{C}\rightarrow\mathbb{R}$ is then given by
@@
Note that the $N$ scalar fields $\phi^{a}$ may be collected into an $N$-tuplet, making manifest the global $O(N)$ symmetry of (61). For the purpose of facilitating the reduction process, we shall make a change of coordinates, such that this $O(N)$ is no longer manifest; in particular, we shall adopt $N$-dimensional polar coordinates $(R,X^{1},\,\cdots\,,X^{N-1})$, defined according to
$\begin{split}\phi^{1}=R\cos X^{1}\\ \phi^{2}=R\sin X^{1}\cos X^{2}\\ \vdots\quad\quad\vdots\quad\quad\vdots\quad\quad\\ \phi^{N-1}=R\sin X^{1}\cdots\,\sin X^{N-2}\cos X^{N-1}\\ \phi^{N}=R\sin X^{1}\cdots\,\sin X^{N-2}\sin X^{N-1}\end{split}$ (62)
Writing the radial coordinate $R$ as $e^{\rho/2}$, the Lagrangian (61) may be expressed as
$L=e^{\rho}\sqrt{-g}\left[g^{\mu\nu}\left(\frac{1}{8}\rho_{\mu}\rho_{\nu}+\sum_{a=1}^{N-1}\frac{1}{2}X_{\mu}^{a}X_{\nu}^{a}\right)-\frac{1}{2}m^{2}\right]$ (63)
The Poincaré-Cartan form $\Theta_{L}$ is simply
$\Theta_{L}=e^{\rho}\sqrt{-g}\,g^{\mu\nu}\left(\frac{1}{4}\rho_{\nu}\,d\rho+\sum_{a=1}^{N-1}X^{a}_{\nu}\,dX^{a}\right)\wedge\,d^{3}x_{\mu}-e^{\rho}\sqrt{-g}\left[g^{\mu\nu}\left(\frac{1}{8}\rho_{\mu}\rho_{\nu}+\sum_{a=1}^{N-1}\frac{1}{2}X_{\mu}^{a}X_{\nu}^{a}\right)+\frac{1}{2}m^{2}\right]d^{4}x$
from which, in accordance with the procedure outlined in section (IV.1), we compute the 1-forms $\theta_{L}^{\mu}$
$\theta_{L}^{\mu}=e^{\rho}\sqrt{-g}\,g^{\mu\nu}\left(\frac{1}{4}\rho_{\nu}\,d\rho+\sum_{a=1}^{N-1}X^{a}_{\nu}\,dX^{a}\right)$
@@
Finally, we eliminate the derivative coordinate $\rho_{\mu}$ in favour of the action density, in order to obtain the following Herglotz Lagrangian
$L^{H}=\sqrt{-g}\left[\frac{1}{2}\sum_{a=1}^{N-1}g^{\mu\nu}X^{a}_{\mu}X^{a}_{\nu}-\frac{1}{2}m^{2}\right]-\frac{2}{\sqrt{-g}}g_{\mu\nu}s^{\mu}s^{\nu}$ (64)
Geometrically speaking, this Herglotz Lagrangian is a function on the space
@@
in which $\widetilde{E}\cong M\times\mathbb{R}^{N-1}$ is a configuration space formed by taking the quotient of $E=M\times\mathbb{R}^{N}$ by the one-dimensional orbits of $\Sigma=\partial_{\rho}$. Additionally, $L^{H}$ is insensitive to internal rotations amongst the $X^{a}$, and so the reduced theory possesses a global O(N1)O(N-1) symmetry.
We have written the Herglotz Lagrangian in a way which makes manifest the elegant physical interpretation of our reduction procedure. We began with a conservative theory of $N$ massive, non-interacting scalar fields; this system possessed a scaling symmetry, in which the superfluous degree of freedom could be considered to be a global radial scale. Upon eliminating this variable, we obtain a theory of (N1)(N-1) massless scalar fields subject to a constant potential, whose strength is fixed by the original mass, together with a friction-like term which arises due to the inaccessibility of the radial scale. Moreover, from the form of the Lagrangian, it is apparent that the massless scalar dynamics appear completely decoupled from the friction-like piece.
One may be inclined to object that the two components are in fact coupled, due to the presence of the factor of $\textrm{Sym}_{2}^{d-1,1}(M)$. However, this not the case, as the background is parametric, and so $\textrm{Sym}_{2}^{d-1,1}(M)$ simply describes the space of all possible choices of metric, and is in no way dynamical. In light of this, we may write the Herglotz Lagrangian (64) as the sum of two independent pieces: $L^{H}=L^{H}_{N-1}+L^{H}_{\textrm{fric}}$, in which $L^{H}_{N-1}$ is the Lagrangian of the (N1)(N-1) massless scalar fields on $J^{1}\widetilde{E}\times_{M}\textrm{Sym}^{d-1,1}_{2}(M)$, and $L^{H}_{\textrm{fric}}$ refers to the frictional component, parameterised by the action density.
This separation into a Lagrangian on a lower-dimensional jet bundle and an action-dependent piece is highly attractive from the perspective of quantisation. Unfortunately, it is far from clear how such a process should proceed, and, as we shall see in the second example, the decoupling of the frictional component from the reduced space dynamics does not survive within an interacting field theory.
Let us now examine the corresponding contact Hamiltonian; the most straightforward way to obtain $H^{c}$ would be to perform a Legendre transform on (64); however, for didactical purposes, we shall compute the full Hamiltonian from (63), and carry out the reduction procedure using the methods of section (IV.2).
The field momenta $p^{\mu}_{\rho}$ and $p^{\mu}_{a}$ corresponding to $\rho$ and $X^{a}$ respectively are
$p^{\mu}_{\rho}=\frac{\partial L}{\partial\rho_{\mu}}=\frac{1}{4}e^{\rho}\sqrt{-g}\,g^{\mu\nu}\rho_{\nu}\quad\quad\quad\quad\quad p^{\mu}_{a}=\frac{\partial L}{\partial X^{a}_{\mu}}=e^{\rho}\sqrt{-g}\sum_{a=1}^{N-1}g^{\mu\nu}X_{\nu}^{a}$ (66)
@@
$H=\frac{1}{e^{\rho}\sqrt{-g}}g_{\mu\nu}\left(2p^{\mu}_{\rho}p^{\nu}_{\rho}+\frac{1}{2}\sum_{a=1}^{N-1}p_{a}^{\mu}p_{a}^{\nu}\right)+\frac{1}{2}e^{\rho}\sqrt{-g}\,m^{2}$ (67)
@@
$\omega_{h}^{\mu}=d\rho\wedge dp^{\mu}_{\rho}+\sum_{a=1}^{N-1}dX^{a}\wedge dp_{a}^{\mu}$
@@
$D=\frac{\partial}{\partial\rho}+p^{\mu}_{\rho}\frac{\partial}{\partial p^{\mu}_{\rho}}+\sum_{a=1}^{N-1}p^{\mu}_{a}\frac{\partial}{\partial p^{\mu}_{a}}$ (68)
@@
$s^{\mu}=\frac{p^{\mu}_{\rho}}{e^{\rho}}\quad\quad\quad\quad\quad\Pi^{\mu}_{a}=\frac{p^{\mu}_{a}}{e^{\rho}}$ (69)
@@
$H^{c}=\frac{H}{e^{\rho}}=\frac{1}{2\sqrt{-g}}\sum_{a=1}^{N-1}g_{\mu\nu}\Pi^{\mu}_{a}\Pi^{\nu}_{a}+\frac{1}{2}\sqrt{-g}\,m^{2}+\frac{2}{\sqrt{-g}}g_{\mu\nu}s^{\mu}s^{\nu}\quad\quad\quad\quad\eta^{\mu}=\frac{\iota_{D}\omega^{\mu}_{h}}{e^{\rho}}=ds^{\mu}-\sum_{a=1}^{N-1}\Pi^{\mu}_{a}\,dX^{a}$ (70)
The multiphase space upon which the original Hamiltonian (67) was defined is $J^{1}E^{*}\times_{M}\textrm{Sym}_{2}^{d-1,1}(M)$, in which we recall that $J^{1}E^{*}$ is the multimomentum bundle over $E=M\times\mathbb{R}^{N}$. Having made the contact reduction, $H^{c}$ is now a function on
@@
where, as in (65), $\widetilde{E}=M\times\mathbb{R}^{N-1}$. Unsurprisingly, as in the Lagrangian formalism, we observe that our contact Hamiltonian (70) consists of a piece describing a theory of (N1)(N-1) massless scalar fields, with a constant shift in the energy, together with an action-dependent term, whose dissipative nature serves to compensate the inaccessibility of the global scale variable.
@@
In the previous example, we considered the case of $N$ non-interacting massive scalar fields, and observed that, dynamically, such a system is entirely equivalent to a configuration of (N1)(N-1) non-interacting massless scalar fields, together with a decoupled friction-like component. Any field theory of physical interest necessarily contains interactions amongst its components; for this reason, we shall now analyse a simple model of two real, interacting scalar fields $\phi$ and $\chi$.
Working over the two-dimensional spacetime manifold $M=\mathbb{R}^{1,1}$, with coordinates $(t,x)$, and fixed Minkowski metric $\eta_{\mu\nu}=\textrm{diag}(1,-1)$, we have the covariant configuration space $E=M\times\mathbb{R}^{2}$. The first jet bundle $J^{1}E$ admits local coordinates $(t,x,\phi,\chi,\phi_{\mu},\chi_{\mu})$, and the Lagrangian function we shall consider is
@@
Here, since we have only two spacetime coordinates, we have chosen to label the velocities as $\phi_{t}$, rather than $\phi_{0}$, for example. Evidently, this Lagrangian is highly non-physical, and should therefore be considered to be nothing more than a convenient and illustrative example. Additionally, there are a number of dimensionful coupling constants, which, for simplicity of exposition, have been set to unity. In the previous example, significant emphasis was placed on the geometry of the symmetry-reduced space, and subsequent physical interpretation of the final theory; in order to ensure an adequate presentation of all facets of the formalism we have developed throughout, we shall here strive to illustrate the process by which multivector field solutions to the equations of motion are obtained.
It is relatively easy to see from (72) that $\phi$ is our scaling variable; indeed, in keeping with the notation developed throughout section (IV.1), we have
$\bar{\Sigma}=\frac{1}{2}\phi\frac{\partial}{\partial\phi}\quad\quad\implies\quad\quad\Sigma=\frac{1}{2}\left(\phi\frac{\partial}{\partial\phi}+\phi_{t}\frac{\partial}{\partial\phi_{t}}+\phi_{x}\frac{\partial}{\partial\phi_{x}}\right)$
which is a scaling symmetry of degree one. Adopting coordinates along the scaling direction, we change field variables according to $\phi\mapsto e^{\rho/2}$. Consequently, the Lagrangian (72) may be expressed as
@@
$\begin{split}\Theta_{L}=e^{\rho}\biggr[\left(\frac{1}{4}\rho_{t}+\frac{1}{2}\left(\chi_{t}-\chi_{x}\right)\right)\,d\rho\wedge dx+\left(\frac{1}{4}\rho_{x}+\frac{1}{2}\left(\chi_{t}-\chi_{x}\right)\right)\,d\rho\wedge dt+\left(\chi_{t}+\frac{1}{2}\left(\rho_{t}-\rho_{x}\right)\right)\,d\chi\wedge dx\\ +\,\left(\chi_{x}+\frac{1}{2}\left(\rho_{t}-\rho_{x}\right)\right)\,d\chi\wedge dt-\left(\frac{1}{8}(\rho_{t}^{2}-\rho_{x}^{2})+\frac{1}{2}(\chi_{t}^{2}-\chi_{x}^{2})+\frac{1}{2}(\rho_{t}-\rho_{x})(\chi_{t}-\chi_{x})+\chi\right)dt\wedge dx\biggr]\end{split}$ (74)
We shall now conduct an analysis of the dynamics described by this multisymplectic Lagrangian system, before carrying out the symmetry reduction. We then conclude this example, comparing the on-shell behaviour of each of the two theories. A straightforward computation of the Euler-Lagrange field equations for $\rho$ yields
@@
In order to illustrate how these equations of motion may be deduced from a more geometrical perspective, we shall assume that there exist locally decomposable $\widehat{\pi}-$transverse multivector field solutions $\boldsymbol{X}_{L}\in\textrm{ker}_{\widehat{\pi}}^{2}\,\Omega_{L}$, and write
@@
The equations of motion could then be recovered geometrically, via an analysis similar to that carried out above; recall, however, that the multicontact 3-form $\Omega_{L^{H}}$ is constructed from the twisted differential operator $\Omega_{L^{H}}=\bar{d}\Theta_{L^{H}}$, as defined in (27). We shall not conduct such an analysis here, since the procedure is entirely analogous to that which we have already illustrated, and the results of the geometric formulation are known to be equivalent to the content of the Herglotz-Lagrange field equations.
The elimination of the scaling coordinate from this theory has an interesting physical interpretation. In the previous example, we considered the case of non-interacting scalar fields, where it was observed that the Herglotz Lagrangian assumed the form of a sum of a piece describing the reduced space scalar field dynamics, together with an independent component containing the action density variables. In this case, by contrast, when the original scalar fields exhibit some kind of interaction (which here, is mediated by the mixed derivative term $(\rho_{t}-\rho_{x})(\chi_{t}-\chi_{x})$) we observe that the corresponding Herglotz Lagrangian contains terms which couple $\chi$ to the action density. In this way, upon eliminating the scaling variable $\rho$, the coupling between the two original scalar fields manifests as an interaction between the remaining field and the frictional degrees of freedom.

### v4

@@ VII Field Theory Classification
All theories we shall consider will possess a metric; however, the precise status of this metric within the theory of interest will affect the underlying multisymplectic structure. In the simplest of cases, the metric is a fixed object, so that it neither evolves dynamically, nor does it assume different values. This occurs most often when considering theories on a flat static background, so that $g_{\mu\nu}=\eta_{\mu\nu}$. Another possibility is that the metric enters the theory parametrically; in this case, we may choose to prescribe a particular background, as the context befits, but once chosen, it is fixed. For this reason, the components of the metric are included as additional variables, while the corresponding velocities are not. In practice, when a theory possesses a metric of this type, we modify the configuration space and jet bundle as follows
@@
in which $\textrm{Sym}_{2}^{d-1,1}(M)$ denotes the space of symmetric covariant tensors of rank two and Lorentzian signature (d1,1)(d-1,1) on $M$; sections of $\textrm{Sym}_{2}^{d-1,1}(M)$ correspond to metrics on $M$. We do not consider the jet bundle over $E\times_{M}\textrm{Sym}_{2}^{d-1,1}(M)$, but append the factor of $\textrm{Sym}_{2}^{d-1,1}(M)$ to $J^{1}E$; this is precisely so that the components $g_{\mu\nu}$ of the metric are amongst the field variables, but the derivatives $g_{\mu\nu,\rho}$ are not.
The final possibility is that of a variational metric, which typically arises in theories coupled to gravity. As an example, consider a principal $G$-bundle $\pi:P\rightarrow M$, with gauge group $G=U(1)$; denote by $C(P)\rightarrow M$ the bundle of connections of $P$, where we may identify $C(P)\cong J^{1}P/G$ [48]. Coupling this $U(1)$ gauge theory to gravity requires us to work with the following covariant configuration space
@@
upon which $(x^{\mu},A_{\mu};g_{\mu\nu})$ provide local coordinates. For concreteness, we have restricted ourselves to $d=3+1$ spacetime dimensions. The gauge-invariant Lagrangian is now a function on
@@
The use of the jet bundle $J^{1}(\textrm{Sym}_{2}^{3,1}(M))$ introduces velocity variables for the components $g_{\mu\nu}$, and so local coordinates on $\mathcal{P}$ are given by $(x^{\mu},A_{\mu},g_{\mu\nu},A_{\mu,\nu},g_{\mu\nu,\rho})$; when evaluated on the jet prolongation $j^{1}\phi:M\rightarrow\mathcal{P}$ of some section $\phi\in\Gamma(M,E)$, the velocity coordinates become the familiar derivatives $\partial_{\nu}A_{\mu}$ and $\partial_{\rho}g_{\mu\nu}$.
@@
As an illustration of the formalism we have developed, we consider an example of $N$ non-interacting real scalar fields $\phi^{a}$ (with $a=1,\,\cdots,N$) moving in a curved, $d$-dimensional spacetime $M$, equipped with a Lorentzian metric $g_{\mu\nu}$ of signature $(+,-,-,\cdots\,)$, taken to be parametric in nature. In accordance with the discussion above, this requires us to make the replacements
@@
where $E$ is the trivial bundle $E=M\times\mathbb{R}^{N}\rightarrow M$, which would be an appropriate covariant configuration space for a flat background. Local coordinates on $\mathcal{C}$ are $(x^{\mu},\phi^{a},\phi^{a}_{\mu};g_{\mu\nu})$, where we have used a semicolon to separate the variational and parametric degrees of freedom. The Lagrangian function $L:\mathcal{C}\rightarrow\mathbb{R}$ is given by
@@
Let $X^{i}$ $(i=1,\,\cdots,N-1)$ denote local coordinates on $S^{N-1}$, and $n^{a}$ those of the embedding $n:S^{N-1}\hookrightarrow\mathbb{R}^{N}$. The induced metric on $S^{N-1}$, denoted $G_{ij}$, is the pullback of the flat metric $\delta_{ab}$ via the embedding $n$. Introducing a radial coordinate $R:=e^{\rho/2}\in\mathbb{R}^{\times}$, we write $\phi^{a}=e^{\rho/2}n^{a}(X)$, and the Lagrangian (40) may be expressed as
$L=e^{\rho}\sqrt{-g}\left[g^{\mu\nu}\left(\frac{1}{8}\rho_{\mu}\rho_{\nu}+\frac{1}{2}G_{ij}X_{\mu}^{i}X_{\nu}^{j}\right)-\frac{1}{2}m^{2}\right]$ (41)
From this Lagrangian, and making use of (3), we compute the 1-forms $\theta_{L}^{\mu}$
$\theta_{L}^{\mu}=e^{\rho}\sqrt{-g}\,g^{\mu\nu}\left(\frac{1}{4}\rho_{\nu}\,d\rho+G_{ij}X^{j}_{\nu}\,dX^{i}\right)$
@@
Writing $L^{H}=f-\rho_{\mu}s^{\mu}$, and eliminating the velocity coordinate $\rho_{\mu}$, we find
$L^{H}=\sqrt{-g}\left[\frac{1}{2}g^{\mu\nu}G_{ij}X^{i}_{\mu}X^{j}_{\nu}-\frac{1}{2}m^{2}\right]-\frac{2}{\sqrt{-g}}g_{\mu\nu}s^{\mu}s^{\nu}$ (42)
Geometrically, upon introducing polar coordinates, we see that $E=(M\times\mathbb{R}^{N})\backslash\{R=0\}\cong M\times(\mathbb{R}_{+}\times S^{N-1})$. The elimination of the scaling degree of freedom corresponds to forming the quotient space under the one-dimensional orbits of $\Sigma$. We denote this space $\widetilde{E}$, and deduce that $\widetilde{E}\cong M\times S^{N-1}$. Consequently, the Herglotz Lagrangian may be identified as a function on the space
@@
From a conservative theory of $N$ massive, non-interacting scalar fields, a change of coordinates allowed us to identify a superfluous radial scaling variable. Upon eliminating this variable, we obtain a theory of (N1)(N-1) massless scalar fields subject to a constant potential, whose strength is fixed by the original mass. Further, as a consequence of the change of variables, the massless scalar fields acquire a non-trivial internal metric. Finally, the reduced theory possesses no notion of global scale, and this is compensated by an action-dependent term, which is frictional in nature. From the form of the Lagrangian, it is apparent that the massless scalar dynamics appear completely decoupled from the friction-like piece, allowing us to express (42) as the sum $L^{H}=L^{H}_{N-1}+L^{H}_{\textrm{fric}}$, in which $L^{H}_{N-1}$ is the Lagrangian of the (N1)(N-1) massless scalar fields on $J^{1}\widetilde{E}\times_{M}\textrm{Sym}^{d-1,1}_{2}(M)$, and $L^{H}_{\textrm{fric}}$ refers to the frictional component, parameterised by the action density.
Let us now examine the corresponding contact Hamiltonian; the most expedient way to obtain $H^{c}$ would be to perform a Legendre transform on (42); however, for didactical purposes, we shall compute the full Hamiltonian from (41), and carry out the reduction using the methods of section (VI). The field momenta $p^{\mu}_{\rho}$ and $p^{\mu}_{i}$ corresponding to $\rho$ and $X^{i}$ respectively are
$p^{\mu}_{\rho}=\frac{\partial L}{\partial\rho_{\mu}}=\frac{1}{4}e^{\rho}\sqrt{-g}\,g^{\mu\nu}\rho_{\nu}\quad\quad\quad\quad\quad p^{\mu}_{i}=\frac{\partial L}{\partial X^{i}_{\mu}}=e^{\rho}\sqrt{-g}g^{\mu\nu}G_{ij}X_{\nu}^{j}$ (44)
@@
$H=\frac{1}{e^{\rho}\sqrt{-g}}g_{\mu\nu}\left(2p^{\mu}_{\rho}p^{\nu}_{\rho}+\frac{1}{2}G^{ij}p_{i}^{\mu}p_{j}^{\nu}\right)+\frac{1}{2}e^{\rho}\sqrt{-g}\,m^{2}$ (45)
@@
$\omega_{H}^{\mu}=d\rho\wedge dp^{\mu}_{\rho}+dX^{i}\wedge dp_{i}^{\mu}$
@@
$D=\frac{\partial}{\partial\rho}+p^{\mu}_{\rho}\frac{\partial}{\partial p^{\mu}_{\rho}}+p^{\mu}_{i}\frac{\partial}{\partial p^{\mu}_{i}}$ (46)
@@
$s^{\mu}=\frac{p^{\mu}_{\rho}}{e^{\rho}}\quad\quad\quad\quad\quad\Pi^{\mu}_{i}=\frac{p^{\mu}_{i}}{e^{\rho}}$ (47)
@@
$H^{c}=\frac{1}{2\sqrt{-g}}g_{\mu\nu}G^{ij}\Pi^{\mu}_{i}\Pi^{\nu}_{j}+\frac{1}{2}\sqrt{-g}\,m^{2}+\frac{2}{\sqrt{-g}}g_{\mu\nu}s^{\mu}s^{\nu}\quad\quad\quad\quad\eta^{\mu}=ds^{\mu}-\Pi^{\mu}_{i}\,dX^{i}$ (48)
The multiphase space upon which the original Hamiltonian (45) was defined is $J^{1}E^{*}\times_{M}\textrm{Sym}_{2}^{d-1,1}(M)$, in which we recall that $J^{1}E^{*}$ is the multimomentum bundle over $E\cong M\times(\mathbb{R}_{+}\times S^{N-1})$. Having made the contact reduction, $H^{c}$ is now a function on
@@
where, as in (43), $\widetilde{E}\cong M\times S^{N-1}$. Unsurprisingly, as in the Lagrangian formalism, we observe that our contact Hamiltonian (48) consists of a piece describing a theory of (N1)(N-1) massless scalar fields, with a constant shift in the energy, together with an action-dependent term, whose dissipative nature serves to compensate the inaccessibility of the global scale variable.
@@
Any field theory of physical interest necessarily contains interactions amongst its components; for this reason, we shall now consider a simple model of two real, interacting scalar fields $\phi$ and $\chi$, and focus our analysis more on the use of multivector fields, so as to ensure an adequate presentation of all facets of the formalism we have developed. Working over the two-dimensional spacetime manifold $M=\mathbb{R}^{1,1}$, with coordinates $(t,x)$, and fixed Minkowski metric $\eta_{\mu\nu}=\textrm{diag}(1,-1)$, we have the covariant configuration space $E=M\times\mathbb{R}_{+}^{2}$. The first jet bundle $J^{1}E$ admits local coordinates $(t,x,\phi,\chi,\phi_{\mu},\chi_{\mu})$, and the Lagrangian function we shall consider is
@@
Evidently, this Lagrangian is highly non-physical, and should therefore be considered to be nothing more than a convenient and illustrative example. Additionally, there are a number of dimensionful coupling constants, which, for simplicity, have been set to unity. It is relatively easy to see from (50) that the vector field
$\Sigma=\frac{1}{2}\left(\phi\frac{\partial}{\partial\phi}+\phi_{t}\frac{\partial}{\partial\phi_{t}}+\phi_{x}\frac{\partial}{\partial\phi_{x}}\right)$
is a scaling symmetry of degree one. Adopting coordinates along this direction, we write $\phi\mapsto e^{\rho/2}$, with which the Lagrangian becomes
@@
$\begin{split}\Theta_{L}=e^{\rho}\biggr[\left(\frac{1}{4}\rho_{t}+\frac{1}{2}\left(\chi_{t}-\chi_{x}\right)\right)\,d\rho\wedge dx+\left(\frac{1}{4}\rho_{x}+\frac{1}{2}\left(\chi_{t}-\chi_{x}\right)\right)\,d\rho\wedge dt+\left(\chi_{t}+\frac{1}{2}\left(\rho_{t}-\rho_{x}\right)\right)\,d\chi\wedge dx\\ +\,\left(\chi_{x}+\frac{1}{2}\left(\rho_{t}-\rho_{x}\right)\right)\,d\chi\wedge dt-\left(\frac{1}{8}(\rho_{t}^{2}-\rho_{x}^{2})+\frac{1}{2}(\chi_{t}^{2}-\chi_{x}^{2})+\frac{1}{2}(\rho_{t}-\rho_{x})(\chi_{t}-\chi_{x})+\chi\right)dt\wedge dx\biggr]\end{split}$
We shall now conduct an analysis of the dynamics described by this Lagrangian system, before carrying out the symmetry reduction. We then conclude by comparing the on-shell behaviour of each of the two theories. A straightforward computation of the Euler-Lagrange field equations for $\rho$ yields
@@
In order to illustrate how these equations of motion are deduced from a geometrical perspective, we shall assume that there exist locally decomposable $\widehat{\pi}-$transverse multivector field solutions $\boldsymbol{X}_{L}\in\textrm{ker}_{\widehat{\pi}}^{2}\,\Omega_{L}$, and write
@@
The equations of motion could then be recovered geometrically; however, since this analysis is analogous to that which we have already performed, and the results of the geometric formulation are known to be equivalent to the content of the Herglotz-Lagrange field equations, we shall not pursue this further.
In contrast to the previous example, the presence of the interaction term $(\rho_{t}-\rho_{x})(\chi_{t}-\chi_{x})$ means that the reduced theory does not decouple into a piece describing the scalar dynamics, and another, independent component, containing the action density. Here, the Herglotz Lagrangian contains terms which couple $\chi$ directly to the action density, and so upon eliminating the scaling variable, the coupling between the original scalar fields manifests as an interaction between the remaining field and the frictional degrees of freedom.